# Okapi: Efficiently Safeguarding Speculative Data Accesses in Sandboxed Environments

### v2

@@ 7-C Vaulting
salsa20_stream 287.93 % 376 28.88 % 6
salsa20_stream_xor 171.66 % 440 7.29 % 6
secretbox_easy 170.69 % 1450 16.00 % 27
secretbox_open 182.16 % 1859 15.26 % 29
aead_ietf_encrypt 106.56 % 784 36.38 % 35
aead_ietf_decrypt 103.99 % 784 34.50 % 37
@@
box_easy 154.54 % 214534 18.16 % 1324
box_open 155.04 % 214030 18.33 % 1313
Mean 168.43 % 45244.6 17.86 % 292
We benchmark the Libsodium [17] constant-time cryptographic library which already comes with a page-aligned memory allocator. This feature facilitates achieving poisoning security (cf. Sec. 6-B).
In this experiment we compare two different choices of granularity for $p_{H}$. First, we instrument the hashing, secret and public key routines within the cryptographic library to execute an OkapiReset instruction after every secret-dependent load instruction to prevent poisoning attacks from leaking secret data. This is the most conservative choice of $p_{H}$, also referred to as atomic $p_{H}$. Secondly, we adapt the same routines to reset the permissions on returns. This instrumentation approximates the security-critical parts of the program to the set of sections $p_{H}^{\prime}$ consisting of the code page(s) of entire functions. While this measure may not guarantee full security, it drastically reduces the attack surface as well as the potential gadget space in proportion to the region size (cf. Sec. 9-B). This can be automated w.r.t. to a predefined logical granularity. To this end, we modified the LLVM compiler framework [30]. LLVM’s intermediate representation allows for an easy adaptation of the chosen granularity. We refer to this instrumentation as function-sized $p_{H}^{\prime}$. The results in Tab. II serve as an estimate of the performance gain for the case that the developer can ensure the absence of gadgets in function-sized granularity, thus achieving full poisoning security.
@@
Tab. II summarizes the runtime overhead of the median execution times of 100 runs of the different routines available in the library normalized to the unmodified version on an unmodified CPU. As can be seen, the finest possible granularity of $p_{H}$ introduces an average overhead of 168.43 %. Resorting to the more coarse-grained function-granular choice of $p_{H}^{\prime}$ significantly reduces the incurred overhead down to 17.86 %. Note that the Okapi security features against poisoning complement the security level provided by the constant-time library. Without Okapi, a poisoning attack can redirect the control flow such that the vault leaks the secret despite constant-time measures in place.
8 Hardware Overhead
We implemented the Okapi architecture according to Sec. 6 in the open-source BOOM processor. We ran FPGA synthesis for the mega configuration of BOOM that matches the performance of Intel Core 2 Duo in terms of CoreMark/MHz. In comparison to the baseline design, Okapi increases the number of lookup tables and flip-flops by 0.39 % and 0.35 %, respectively. This confirms our claim that the Okapi architecture incurs a negligible hardware overhead.
9 Security Evaluation
9-A Formal Verification
We base our verification on the methodology of Unique Program Execution Checking (UPEC), as presented in [19], which formally establishes weak speculative non-interference for a design under verification (DUV). Our DUV is BOOM extended by the described Okapi architecture. We adapt UPEC to sandboxing (cf. Sec. 3) by constraining the UPEC property [19] such that the program in its sequential semantics never accesses high memory locations.
The Okapi-adapted UPEC property holds after a proof runtime of 41 minutes. This guarantees breakout security for sandboxes of custom size, provided that the software executes an OkapiReset instruction on sandbox exits and that it places confidential and public data onto different data pages. In addition, this proves that any program $p$ is poisoning-secure if $p_{H}$ does not add the high memory locations to the trust domain.
9-B Attack Surface Analysis
As explained in Sec. 4-A, the number of speculatively accessible pages of a trust domain increases over time. However, the empirical results for SPEC CPU2017 shown in Fig. 2 demonstrate that programs access only 5.69 % of their mapped code pages and 31.27 % of their mapped data pages between two context switches [^2] Every context switch involves a privilege switch, on average. Thus, at most 31.27 % of a program’s data is accessible speculatively by any Spectre read gadget in the mapped address space. As Okapi prevents the execution of gadgets following transient jumps to arbitrary code pages, attackers could only exploit gadgets on the 5.69 % of the code pages that are reached by the program flow non-speculatively. This means that developers only need to care about Spectre gadgets in the immediate vicinity of locations (i.e., the pages) where the code works with secret data. Spectre gadgets in other locations are automatically suppressed by Okapi without any further action needed by the developers or the compiler.
An interesting observation we make is that, even if an attacker could successfully access confidential data, the cache side channel can only be exploited if the address used to encode the data is also within the trust domain. Although it is difficult to quantify the effect of this restriction on poisoning attacks, it is apparent that, without any software support, the Okapi hardware alone increases the necessary effort to exploit poisoning attacks drastically.
10 Related Work
SafeBet [23] exploits temporal locality based on similar observations as in Okapi. However, the SafeBet mitigation scheme introduces a cache-like hardware structure to keep track of which code regions have already accessed which memory regions legally. The performance overhead incurred by SafeBet depends on the chosen granularity of the monitored code regions, and ranges from 6 % (at 1 GB granularity) to over 270 % (at address level). Also, the hardware area overhead is significantly higher than that of Okapi, due to the additional 8.3 KB high-speed cache needed for access mapping.
Hardware-only mitigations based on DIFT, such as STT [48] or SecureBOOM [27], protect all data in the memory and caches. These schemes allow speculative accesses to data and track their effects through the pipeline with taint bits. Instructions capable of leaking their operands (“transmitters”) are then delayed until they are no longer speculative. The performance overheads of these mitigations range from 14.8 % to 36.0 %. DOLMA [31] and SPT [13] extend hardware-only mitigations to also protect the content of the register file (i.e., data that has been accessed non-speculatively) from being leaked speculatively while incurring a performance overhead of up to 45.0 %. Although these techniques (STT [48], SecureBOOM [27], DOLMA [31] and SPT [13]) deliver stronger security guarantees compared to the bare Okapi hardware, they come at the price of higher hardware and performance overheads. They significantly increase the design complexity of the microarchitecture, particularly across the memory hierarchy, which impedes their adoption in existing microarchitectures.
SpectreGuard [20] allows the programmer to annotate sensitive memory regions manually to protect data from Spectre-PHT attacks. The annotations are then used by the hardware to prevent memory accesses under control-flow speculation for all tagged regions by delaying their execution until they are safe to be executed. Protecting the content of the heap comes with 8.0 % performance overhead, while protecting the entire memory space used in the SPEC CPU2006 benchmark suite slows down the system by 20.0 %.
ConTeXT [39] pursues a related approach, where memory regions are marked as non-transient at a page-level granularity by the programmer, preventing access during transient execution. Additionally, ConTeXT employs DIFT to track sensitive data that has been loaded non-transiently into the register file, safeguarding such data from being used in transient operations. It requires intrusive changes over all layers (HW, OS, compilers) and incurs a baseline overhead of 1.26 % without protecting any data. With protection in place, the overhead reaches 71.14 % in security-critical applications.
ProSpeCT [16] also leverages the concept of manual annotations in software to extend secure speculation to constant-time programming, resulting in 10.0 % to 40.0 % performance overhead on custom benchmarks. Similar to ConTeXT, ProSpeCT is capable of protecting data at rest as well as data in the registers from TES attacks.
Okapi effectively enforces dynamic sandboxing for speculative data accesses and therefore shares this objective with conventional sandboxing approaches and (in-)process isolation techniques that also mitigate certain Spectre attacks [34, 38, 35, 7]. However, these mitigations pursue a different security target and introduce a much higher overhead in implementation and performance. In addition, while most of these techniques can thwart Spectre-PHT attacks, many of them leave an open attack surface to Spectre-BTB.
11 Conclusion
At its core, Okapi prevents these attacks by delaying speculative accesses to data pages which have not yet been accessed from within the currently running sandbox. Okapi provides the possibility to customize the granularity of sandboxes, thereby covering a range of possible applications. The Okapi hardware architecture is easy to implement and can be seamlessly integrated into existing designs. In particular, breakout security being highly relevant in applications of edge computing is achieved with an exceptionally low hardware and performance overhead. This bears promise for Okapi’s adoption in industrial practice, possibly replacing software patches deemed too costly.
An interesting direction for further research could be the combination of the Okapi hardware mechanisms with orthogonal software protection mechanisms, e.g., Serberus [33] for further performance improvements.
Acknowledgements
The reported research was partly supported by DFG SPP Nano Security, by BMBF grant SILKOSTU and by the Intel Corp. Scalable Assurance Program.

### v3

@@ Appendix A Additional experimental Results
salsa20_stream 186 186 6
salsa20_stream_xor 194 194 6
secretbox_easy 670 670 27
secretbox_open 808 808 29
aead_ietf_encrypt 337 337 35
aead_ietf_decrypt 337 337 37
@@
box_easy 41,034 41,034 1,324
box_open 41,074 41,074 1,313
Mean 12,576 12,576 292
[…]
@@
This figure summarizes the runtime overhead of the median execution times of 100 runs of the different routines available in the Libsodium library normalized to the unmodified version on an unmodified CPU. As can be seen, the finest possible granularity of $p_{H}$ implemented by OkapiReset introduces an average performance overhead of 168.43 %. While maintaining the same level of security this overhead can be reduced to 59.78 % by using OkapiLoad. Resorting to the more coarse-grained function-granular choice of $p_{H}^{\prime}$ significantly reduces the incurred overhead even further down to 11.60 %.
Figure 11. Libsodium overheads
Appendix B Spectre-PHT Sample Attack
We mounted a Spectre-PHT sample attack from the Google SafeSide repository (Google, 2019) in order to validate the security provided by Okapi. The application holds a public data array as well as a secret data array. We modified the code such that it places the secret data on a separate page. After checking that the attack works as expected on the unmodified out-of-order model, the CPU model is changed to an Okapi out-of-order CPU. The in-process PHT attack performs legal accesses to the public data array to mistrain the predictor until it changes the offset to point to a byte in the secret string. The secret byte is then retrieved by a cache leakage gadget measuring the latency for the entries in an array data structure. Fig. 12 shows the average access time to the elements of that array over ten runs. The unmodified out-of-order CPU model has an average access time of 64 cycles for the 73rd element which represents the character ’I’ indicating a successful attack. For the modified Okapi CPU, on the other hand, no element of the timing array is cached. This can be seen in Fig. 12 where all access latencies are above the threshold, documenting that the Okapi architecture successfully prevents the attack.
[…]
The figure shows the average access latency for the timing array in the Spectre-PHT attack over 10 runs. While the unmodified o3-CPU shows leakage, the Okapi-CPU does not.
Figure 12. Average access latency for the timing array in the Spectre-PHT attack over 10 runs.
Appendix C Synthesis and Place & Route
Synthesis of the BOOMv3 core (Zhao et al., 2020) shows that the front-end dominates the critical path. With both instruction and data cache, as well as parts of the TLB module implemented as SRAM cells, BOOMv3 can easily reach a clock frequency of 500 MHz after synthesis and place & route on a commercial 12 nm FinFET Globalfoundries process for the medium configuration of the core.
We used the worst-case corner as the target corner for the timing optimizations with a supply voltage of 0.72 V slow-slow transistor corner and 125 °C temperature. The employed ARM standard cell library uses only very low threshold transistors for minimum dynamic power and maximum performance. The power results are generated in Synopsys PrimePower with typical corner settings (TT, 0.8 V, 25 °C).
The runtime for the synthesis was roughly 6 hours, with additional 2.5 days for place & route and static timing analysis. Our results for BOOMv3 align well with the reported results for BOOMv2. Here, the developers report a tapeout frequency of 594 MHz with optimized place & route setup in a 28 nm process (Celio, 2017).
Running the same process on our modified OkapiBOOM core shows that the additional hardware does not deteriorate the maximum clock frequency as the implemented design features are not on the critical path. Compared to the baseline design, Okapi adds 0.6 % of cell area. It is interesting to mention that the baseline design takes 1.5 % more area than OkapiBOOM after place & route, due to a more complicated placement and wiring. In addition, incorporating the Okapi features reduces the power consumption by 1.5 %. A detailed comparison is listed in Tab. 3. The small differences are all within the tolerance of the commercial tools (noise), and are not a systematic cause of the Okapi extension.
Our results clearly show that the additional hardware does not jeopardize the synthesizability of the BOOM design. The resulting layout of OkapiBOOM after place & route is shown in Fig. 13.
Table 3. Comparison of Synthesis Results in 12 nm FinFET
Baseline OkapiBOOM
Area*
Combinational Area ($GE$) 859,219 891,426 +3.8 %
Register Area ($GE$) 1,188,868 1,172,435 -1.4 %
Macro Area ($GE$) 554,294 554,294 0.0 %
Total Cell Area ($GE$) 2,602,465 2,618,088 +0.6 %
Total Area ($GE$) 4,730,557 4,659,064 -1.5 %
Power**
Dynamic Power*** ($mW$) 246.7 242.6 -0.6 %
Leakage Power ($mW$) 0.008 0.008 0.0 %
* Gate Equivalents (GE), normalized to single NAND2_X1 size
** at 500 MHz, TT corner (25 °C, VDD = 0.8 V)
***total power = switching power + internal power
[…]
Floorplan of the OkapiBOOM after place & route. It can be seen that the additional Okapi hardware does not jeopardize the synthesizability of the BOOM design.
Figure 13. Floorplan of OkapiBOOM after place & route